# Group Testing with General Correlation Using Hypergraphs

### v1

@@ Abstract
We then generalize the proposed framework of group testing with general correlation in two directions, namely noisy group testing and semi-non-adaptive group testing. In both settings, we provide novel theoretical bounds on the number of tests required.
@@ 1.1 Related Work
Another line of work is quantitative group testing, which it explores scenarios where the test results are not binary but instead, a test returns the exact number of infections within a group. It is shown that in the adaptive setting, $O(\log n)$ tests are enough [Chr80, Aig86] and later the coefficient of $\log n$ was improved[GMSV92]. Recently, [SJ24b] designed low-complexity algorithms with efficient construction and decoding. For non-adaptive group testing, [HKM22] obtains $O(d\log n)$ which matches the information-theoretic lower bounds up to a factor. [SJ24a] has improved the decoding complexity.
@@
In a recent study, [GLS22] studies a scenario in which only specific subsets of individuals can be infected, collectively referred to as the set of candidate subsets denoted as $E$. This model can partly capture the correlation between the nodes. [GLS22] models candidate subsets as hyperedges of a hypergraph defined on the set of individuals (nodes) $V$ and assumes that each hyperedge has a maximum size of $d$. It provides non-adaptive and adaptive group testing strategies that require $O(d\log E)$ and $O(\log|E|+d^{2})$ tests, respectively. These bounds are complemented with the lower bound $\log|E|+d$ for a specific “worst-case” hypergraph. This work is closest to our work, both in terms of the underlying model and the conceptual framework for algorithm design. However, their framework does not allow different likelihoods of the edges. This assumption is not practical for many applications. For instance, very large and very small edges might be unlikelier than the medium-sized edges. Or different classes of individuals may have different risk factors. Notably, our approach encompasses not only the definition of hyperedges but also introduces a distribution over them to introduce likelihood, and as such our work stands in the so called probabilistic group testing paradigm.
@@
There is a variant of group testing, also known as (hyper)graph learning, which is based on graphs and hypergraphs where items are semi-defective, meaning that the test result is positive when specific subsets of the (hyper)graph are included in the test [Aba18, AN19, BS24]. Several bounds have been established for general graphs and hypergraphs [Aba18, AN19], and these bounds have been improved by assuming specific structures on the hypergraphs [BS24]. Although these works utilize the concepts of graphs and hypergraphs, our problem differs in its fundamental nature.
@@
There have been other works where specific structures of graphs and hypergraphs are assumed and the testing is designed by the structure to improve the testing even further. For instance, in [BCN21] and [BS23] they assume that the graph is a line and a block of consecutive nodes is positive. Another example is [NKC+22] where they give theoretical bounds on the number of tests when the graph is a tree or grid. In this work, we do not have any constraints on the structure of the hypergraph.
@@
Probabilistic group testing was studied in [LCH+14], where each individual $i$ is infected with a probability $p_{i}$, independently of others. Their research led to the establishment of near-optimal bounds for both adaptive and non-adaptive group testing, quantified as $O(H(X)+\mu)$, where $\mu=\sum_{i}p_{i}$ represents the average number of infections, and $H(X)=\sum_{i}-p_{i}\log p_{i}$ denotes the entropy. In recent years, there has been a surge of interest in modeling correlations in group testing scenarios. Several noteworthy developments in this area include: modeling correlation using edge-faulty graphs [NKC+22, AU23] where a simple graph $G$ is considered as the underlying contact graph and each edge is dropped with some probability, forming random components. The nodes within a component are then assumed to be in the same state. In [NKC+22], it was shown for several families of graphs that exploiting correlation can lead to significant improvement in the required number of tests (compared to independent group testing). [AU23] shows improvement when the realized graphs have a certain structure, i.e. they are nested. Another line of work is using communities to model correlation [NSG+21a, ACO21, NSG+23, NSG+21b, JCM24]. For example, in [NSG+21a], it is assumed that each community is infected with some probability, and if a community is infected, all of its members are infected with some non-zero probability independent of the others. If a community is not infected, none of its members are infected. For this model, the authors designed efficient adaptive and non-adaptive algorithms that significantly reduce the number of tests by accounting for the structure and pooling the infected communities. They also developed more reliable tests using the structure when the tests are noisy. [ACO21], considers similar communities, but the communities become infected by random initial seeds. The authors design an adaptive algorithm that significantly reduces the number of tests, further supporting the benefits of exploiting correlation.
@@ 1.2 Contributions
  • We model arbitrary statistical correlation by employing hypergraphs, drawing inspiration from [GLS22]. A hypergraph $G=(V,E)$ has node set $V$, $|V|=n$, and edge set $E$, where each hyperedge $e\in E$ is a subset of nodes: $e\subseteq V$. Enhancing the combinatorial model in [GLS22], we consider an arbitrary probability distribution of infection defined over the hyperedges. This captures arbitrary statistical correlation among infection of different nodes.
@@
  • We propose an adaptive group testing algorithm that is capable of exploiting correlation by the means of updating the posterior distribution on the hyperedges given the previous test results. The algorithm works in two stages. Stage 1 focuses on conducting “informative” tests, which significantly narrows down the search space. When these tests are no longer feasible, the algorithm transitions to a second stage where it tests the remaining uncertain nodes individually. We demonstrate that this algorithm successfully identifies the infected set and requires an expected number of tests that can be upper bounded as a function of entropy $H(X)$ and the expected number of infections $\mu$. For distributions that are concentrated around $\mu$, the bound takes the form of ${O}(H(X)+\mu)$.
@@
  • We provide various examples to illustrate the extent of optimality of the algorithm. $H(X)$ is a known lower bound on the number of tests. We show that our proposed algorithm, and the upper bound we establish on the number of tests, ${O}(H(X)+\mu)$, is worst-case order-optimal. We further show that (i) $H(X)$ can be a loose lower bound with a significant gap to the number of tests needed; (ii) for some families of graphs, $\mu$ is a lower bound on the number of tests; and (iii) the lower bound cannot be described using only $\mu$ and $H(X)$. In particular, we provide examples where the tight lower bound is greater than the entropy but less than $\mu$. We further study the performance of the proposed algorithm for random uniform hypergraphs (where each hyperedge of size $d$ is present with probability $r$) and shed light on the optimality of the algorithm as a function of the hypergraph density $r$.
@@
  • We introduce a semi-non-adaptive group testing model, a model in between adaptive and non-adaptive group testing, in which tests are performed sequentially, allowing the algorithm to halt when it finds the infected set. But, the model is semi-non-adaptive in that tests are designed independent of the results of previous tests. For this model, and assuming concentration of the infection size to be around its mean $\mu$, we provide an algorithm that requires ${O}(\mu H(X)+\mu\log n)$ tests. We also highlight challenges in designing non-adaptive algorithms. Specifically, we show that, unlike classic independent group testing, there can be a significant gap in the number of tests needed in adaptive vs non-adaptive testing.
  • We extend our methods and results to noisy group testing (where test results are corrupted by a symmetric noise) in both adaptive and semi-non-adaptive settings.
@@ 2.1 Generalized Group Testing
We further introduce a semi-non-adaptive group testing model (SNAGT) that serves as an intermediate approach between non-adaptive and adaptive group testing. This model is relevant because it combines the advantages of both strategies: the design of individual tests is non-adaptive, meaning each test is planned without relying on the outcomes of previous tests. However, it introduces adaptiveness in the decision to stop testing, allowing the total number of tests, L, to be determined based on prior results. This is particularly practical in group testing scenarios where delays make it challenging to wait for each test result before deciding the next step. The ability to stop testing when necessary provides a flexible and efficient solution without the need for constant adjustments during the testing process. In Section 3.4, we present an algorithm for the adaptive testing. In Section 7, we examine non-adaptive and semi-non-adaptive algorithms, highlighting the challenges associated with each.
@@ 2.3 Preliminaries
The proof provided for the $(1-\epsilon)H(X)$ lower bound is quite general and can be extended to our work as following:
@@
For any algorithm that recovers the sampled edge $e^{*}$ with probability at least $1-\epsilon$ using $T$ tests, we have
@@ 3.3 Overview and Ideas
In Corollary 4.6, we show that in Stage 1, where each test removes at least $c$ fraction of the mass, the algorithm performs at most $O(H(X))$ tests in expectation which is near-optimal. In Stage 2, where high probability nodes are tested individually, we show in Observation 4.7 that the number of tests depends on the expected number of infections at that point, $\tilde{\mu}$. This expectation is obtained by the posterior probability after performing all the previous tests. We prove bounds on $\tilde{\mu}$ under assumptions on the size of the edges. Specifically, when the size of the edges is concentrated, we get a bound in the form of $O(\mu+H(X))$, which is stated in Theorem 4.8. Another way to bound the number of tests in Stage 2 is to first drop the edges that are much larger than $\mu$, which have a low mass, and then run the algorithm. In this case, we can guarantee that $\tilde{\mu}$ is not much larger than $\mu$, and the guarantee for this approach is shown in Theorem 4.10.
@@ 3.4 The Proposed Algorithm
1 Input: Graph $G=(V,E)$, distribution $\mathcal{D}=\mathcal{D}$ over $E$, $c<1/2$, and the set $W$ containing negative nodes.
@@ 5 If $c\leq w(S)\leq 1-c$, test $V\setminus S$:
6     If the test is positive, remove $E(S)$ from $\mathcal{D}$ by def 3.2 to update the posterior distribution.
7     If the test is negative, remove $E\setminus E(S)$ from $\mathcal{D}$ by def 3.2 to update the posterior distribution. Set $W=W\cup(V\setminus S)$.
@@ 9 Else:
14        If the test is positive, remove $E(S)$ from $\mathcal{D}$ by def 3.2 to update the posterior distribution and go to line 1.
15        If the test is negative, test every node $v\in S$ individually with $10<q_{v}<$ 1, update the posterior after every test and return the nodes with $q_{v}=1$.
@@
The proposed algorithm is presented in Algorithm  1. The algorithm undertakes two broad tasks repeatedly
@@
    Determine the set of nodes it tests, denoted as $V\setminus S$, (it identifies a set $S$ whose complement it tests);
@@
We now describe how these broad tasks are executed. In Stage 1 of the algorithm which corresponds to lines 1-1, the algorithm greedily finds a set $S$ so that $c\leq w(S)\leq 1-c$. This is captured in lines 1 and 1. When the set $S$ is found, $V\setminus S$ is tested in line 1 and the posterior is updated accordingly. Such a test helps rule out edges with total weight of at least $c$ are ruled out of contention for being target edge regardless of the outcome of the test. If the algorithm can not find a set $c\leq w(S)\leq 1-c$ at this stage, it tests $V\setminus S$ in line 1 which has a high probability of being negative. If at Line 1, the test result is positive (unexpectedly), the algorithm rules out $1-c$ fraction of the mass and gets back to line 1. Stage 2 of the algorithm is outlined in line 1, where it tests uncertain high probability nodes individually. Note that at this point, all the nodes outside of $S$ are negative.
@@
We now provide details of the sequence followed in the algorithm. At the start of the algorithm, typically the set of nodes that are known to be negative, $W$, is empty. Thus in line 1, the set $S$ is initialized as $V\setminus W=V.$ Clearly in line 1 $w(S)=1$, and $S$ does not satisfy the condition in line 1. The algorithm would now move to line line 1 (typically the condition in line 1 would also fail as the target edge will not be clear at this point). In line 1, the algorithm identifies individual nodes which are tested in the loop starting from line 1, and testing these individual nodes rules out edges with total weight at least $c$ regardless of the outcome of the test. After each test, posteriori probabilities are modified and the graph is modified as in (2) (this modification happens in lines 1,1 and in line 1). The above sequence is repeated until there does not exist individual nodes such that by removing them $S$ satisfies the condition in line 1. Note that $w(S)=1$ each time the algorithm reaches lines 1. The tests are all individual while this duration lasts.
@@
Once the duration described in the previous paragraph ends the algorithm moves to line 1. Note that $w(S)=1$ the first time that the algorithm reaches line 1 under the then posterior probabilities. When the algorithm reaches line line 1, $S$ consists only of nodes that are positive either with very high probability (such that if any one of those is removed, $w(S)$ falls below $c$ from $1$), or with very low probability (such that $w(S)$ exceeds $1-c$ even when any of those is removed from $S$). In line 1, the low probability nodes are identified one at a time, and once such a low probability node is identified, line 1 determines if now there exists a node whose removal ensures that $w(S)$ satisfies the condition in line 1. If there exists such a node, in line 1 it is tested together with the low probability node identified in line line 1. If there does not exist such a node, then the algorithm gets back to line 1 again and removes another low probability node and the same process is repeated. If no node is found in line 1, then all low probability nodes are tested together in line 1. Thus, tests are no longer individual once the algorithm reaches line 1. When a group of nodes is tested, either in line 1 or line 1, the graph is modified as in (2). If the test in line line 1 is negative, then all tested nodes are removed from $V$, and the remaining nodes are all positive with high probability and are tested individually in line line 1. If the test is positive, then only edges consisting of at least $1$ node in the set are retained, and the algorithm returns to line 1 and the sequence is repeated.
Before we start analyzing, we provide some details to get a better intuition of the algorithm and present two example. If an individual node (set) is tested negative, since it is removed from $V$, it is never tested again. If an individual node (set) is tested positive, only edges intersecting that node (set) remain in the graph, thus the weight of the complement of that node (set) becomes 00. Thus, the complement set will never have weight in the interval $[c,1-c]$ and is therefore not tested again.

### v2

@@ Abstract
We then generalize the proposed framework of group testing with general correlation in two directions, namely semi-non-adaptive group testing and noisy group testing. In both settings, we provide novel theoretical bounds on the number of tests required.
@@ 1.1 Related Work
Another line of work is quantitative group testing which explores scenarios where the test results are not binary but instead, a test returns the exact number of infections within a group. It is shown that in the adaptive setting, $O(\log n)$ tests are enough [Chr80, Aig86] and later the coefficient of $\log n$ was improved [GMSV92]. Recently, [SJ24b] designed low-complexity algorithms with efficient constructions and decoding. For non-adaptive group testing, [HKM22] obtains $O(d\log n)$ which matches the information-theoretic lower bounds up to a factor. [SJ24a] has improved the decoding complexity.
@@
In a recent study, [GLS22] studies a scenario in which only specific subsets of individuals can be infected, collectively referred to as the set of candidate subsets denoted as $E$. This model can partly capture correlation between the nodes. [GLS22] models candidate subsets as hyperedges of a hypergraph defined on the set of individuals (nodes) $V$ and assumes that each hyperedge has a maximum size of $d$. It provides non-adaptive and adaptive group testing strategies that require $O(d\log E)$ and $O(\log|E|+d^{2})$ tests, respectively. These bounds are complemented with the lower bound $\log|E|+d$ for a specific “worst-case” hypergraph. This work is closest to our work, both in terms of the underlying model and the conceptual framework for algorithm design. However, their framework does not allow different likelihoods of the edges. This assumption is not practical for many applications. For instance, very large and very small edges might be unlikelier than the medium-sized edges. Or different classes of individuals may have different risk factors. Notably, our approach encompasses not only the definition of hyperedges but also introduces a distribution over them to introduce likelihood, and as such our work stands in the so called probabilistic group testing paradigm.
@@
There is a variant of group testing, also known as (hyper)graph learning, which is based on graphs and hypergraphs where items are semi-defective, meaning that a test result is positive if there is a hyperedge that is completely contained in the test [Aba18, AN19, BS24]. Several bounds have been established for general graphs and hypergraphs [Aba18, AN19], and these bounds have been improved by assuming specific structures on the hypergraphs [BS24]. Although these works utilize the concepts of graphs and hypergraphs, our problem differs in its fundamental nature.
@@
There have been other works where specific classes of graphs or hypergraphs are assumed in the design of group testing. For instance, [BCN21] and [BS23] assume that the graph is a line and a block of consecutive nodes is positive. Another example is [NKC+22] where theoretical bounds are provided on the number of tests when the graph is a tree or grid. In this work, we do not impose any constraints on the structure of the hypergraph.
@@
Probabilistic group testing was studied in [LCH+14], where each individual $i$ is infected with a probability $p_{i}$, independently of others. [LCH+14] establishes near-optimal bounds for both adaptive and non-adaptive group testing, quantified as $O(H(X)+\mu)$, where $\mu=\sum_{i}p_{i}$ represents the average number of infections, and $H(X)=\sum_{i}-p_{i}\log p_{i}$ denotes the entropy. In recent years, there has been a surge of interest in modeling correlations in group testing scenarios. [NKC+22, AU23] model correlation using edge-faulty graphs where a simple graph $G$ is considered as the underlying contact graph and each edge is dropped with some probability, forming random components. The nodes within a component are then assumed to be in the same state. In [NKC+22], it was shown for several families of graphs that exploiting correlation can lead to significant improvement in the required number of tests (compared to independent group testing). [AU23] shows improvement when the realized graphs have a certain structure, i.e. they are nested. Another line of work is using communities to model correlation [NSG+21a, ACO21, NSG+23, NSG+21b, JCM24]. For example, in [NSG+21a], it is assumed that each community is infected with some probability, and if a community is infected, all of its members are infected with some non-zero probability independent of the others. If a community is not infected, none of its members are infected. For this model, the authors designed efficient adaptive and non-adaptive algorithms that significantly reduce the number of tests by accounting for the structure and pooling the infected communities. They also developed more reliable tests using the structure when the tests are noisy. [ACO21], considers similar communities, but the communities become infected by random initial seeds. The authors design an adaptive algorithm that significantly reduces the number of tests, further supporting the benefits of exploiting correlation.
@@ 1.2 Contributions
  • In Section 2, we model arbitrary statistical correlation by employing hypergraphs, drawing inspiration from [GLS22]. A hypergraph $G=(V,E)$ has node set $V$, $|V|=n$, and edge set $E$, where each hyperedge $e\in E$ is a subset of nodes: $e\subseteq V$. Enhancing the combinatorial model in [GLS22], we consider a probability mass function over the hyperedges. This captures arbitrary statistical correlation among infection of different nodes. We design group testing algorithms that return the set of infected nodes with a probability that converges to $1$ as $n$ increases to infinity for each of the scenarios described below.
@@
  • In Section 3, we propose an adaptive group testing algorithm that is capable of exploiting correlation by the means of updating the posterior distribution on the hyperedges given the previous test results. The algorithm works in two stages. Stage 1 focuses on conducting “informative” tests, which significantly narrows down the search space. When these tests are no longer feasible, the algorithm transitions to a second stage where it tests the remaining uncertain nodes individually. We demonstrate that this algorithm successfully identifies the infected set and requires an expected number of tests that can be upper bounded as a function of entropy $H(X)$ and the expected number of infections $\mu$ in Theorem 4.11. We also show that when a stochastic upper bound $u$ on the number of infected nodes is known (ie, the probability that the number of infected nodes is upper bounded by $u$ converges to $1$ as the number of nodes increases to infinity), then $O(H(X)+u)$ tests are sufficient in expectation in Corollary 4.12. Thus, the guarantee on the expected number of tests becomes better as the upper bound becomes tighter. If the number of infections is concentrated around $\mu$, the average number of infections, then $O(\mu)$ constitutes the stochastic upper bound and hence $O(H(X)+\mu)$ tests are sufficient. In Section 5, we apply our algorithm to several existing group testing frameworks and compare our results against these prior approaches.
@@
  • In Section 6, we provide various examples to illustrate the extent of optimality of the algorithm. $H(X)$ is a known lower bound on the number of tests. We show that our proposed algorithm, and the upper bound we establish on the number of tests, ${O}(H(X)+\mu)$, is worst-case order-optimal. We further show that (i) $H(X)$ can be a loose lower bound with a significant gap to the number of tests needed ; (ii) for some families of graphs, $\mu$ is a lower bound on the number of tests; and (iii) the lower bound cannot be described using only $\mu$ and $H(X)$. In particular, we provide examples where the tight lower bound is greater than the entropy but less than $\mu$. We subsequently consider random $d$-regular hypergraphs, where each hyperedge of size $d$ is present with probability $r$. In Section 6.4, we show that the algorithm is order optimal for the expected number of tests in dense $d$-regular random hypergraphs (ie, when $r$ is large), and a modification thereof attains order optimality in sparse $d$-regular random hypergraphs (ie, when $r$ is small); some of the proofs use mild additional conditions that arise naturally.
@@
  • In Section 2, we introduce a semi-non-adaptive group testing model, a model in between adaptive and non-adaptive group testing, in which tests are performed as per a predetermined sequence, which is designed before result of any test is known, but the tests can be terminated any time depending on the results of the already executed tests. In Section 7, we provide an algorithm that requires ${O}(uH(X)+u\log n)$ tests in this model, where $u$ is a stochastic upper bound on the number of infected nodes as defined in the second bullet (Theorem 7.1). The guarantee is clearly worse than that we obtain in a fully adaptive setting. But, this algorithm is easier to execute in that the testing sequence is preplanned and therefore tests can be scheduled in advance. We also highlight challenges in designing non-adaptive algorithms. Specifically, we show that, unlike classic independent group testing, there can be a significant gap in the number of tests needed in adaptive vs non-adaptive testing.
  • In Section 8, we extend our methods and results to noisy group testing (where test results are corrupted by a symmetric noise). The impact of such noise can clearly be mitigated by repeating each test a certain number of times and going with the majority verdict, such repetition however increases the overall number of tests. In Theorem 8.1, we show that in the adaptive scenario the number of repetitions can be substantially reduced by considering the impact of the error in update of the posterior probabilities of edges after each test. The guarantees we obtain on the expected number of overall tests become better as the probability of error reduces.
@@ 2.1 Generalized Group Testing
We further introduce a semi-non-adaptive group testing model (SNAGT) that serves as an intermediate approach between non-adaptive and adaptive group testing. This model is relevant because it combines the advantages of both strategies: the design of individual tests is non-adaptive, meaning each test is planned without relying on the outcomes of previous tests. However, it introduces adaptiveness in the decision to stop testing, allowing the total number of tests, L, to be determined based on prior results. From a practical perspective, the utility of any SNAGT testing strategy is that since who to test at any point in the testing sequence can be determined apriori (ie before any result is known), the individuals who must be tested at any point can be informed well in advance, actually at the start of the overall testing process. This renders the testing process as convenient for the subjects. Cancelation of tests can however be on the fly, as results become available. In Section 3.4, we present an algorithm for the adaptive testing. In Section 7, we examine non-adaptive and semi-non-adaptive algorithms, highlighting the challenges associated with each.
@@ 2.3 Preliminaries
The proof provided for the $(1-\epsilon)H(X)$ lower bound is quite general and can be extended to our work as follows:
@@
For any algorithm that recovers the sampled edge $e^{*}$ with probability at least $1-\epsilon$ using $L$ tests, we have
@@ 3.2 Some Useful Definitions
We further elucidate edges and relation between weights of a set of nodes, probabilities that the nodes in it are infected, probabilities that edges in it are target edges using an additional example. We illustrate the relations that are contrary to common intuition:
  1. 1.
    An edge $e_{1}$ can be a proper subset of another edge, $e_{2}$, but the probability that $e_{1}$ is target edge can be considerably higher than the probability that $e_{2}$ is a target edge.
  2. 2.
    Weight of a set $S$ of nodes is not equal to the sum of the probabilities that the nodes in it are infected, the former can be considerably lower than the latter, the former can be 00 even when it contains several nodes which are infected with probability close to $1$.
Example 3.3.
Consider a network with $k$ islands where each island has $m$ nodes, so $n=mk$. Nodes on the same island are perfectly correlated, ie, they always have the same state, as such every hyper edge always contains all or none of any given island (Figure 3, $k=4,m=3$). Thus, there are $2^{k}$ hyperedges at most; note that in this case the number of hyper edges can be significantly smaller than the size of the power set of nodes, $2^{mk}$. Let nodes in different islands be independent. Consider a setting with three types of islands: low, moderate and high, characterized by their probability of infection. For all nodes in high islands, $p_{i}=1-\gamma$ (e.g. island 4 in Figure 3), where $\gamma$ is a small positive number. In moderate islands (e.g. islands 2,3), $p_{i}\approx\frac{1}{2}$ and in low islands, e.g. island 1, $p_{i}=\epsilon$, where $\epsilon$ is a small, positive number.
Nodes in an edge can be a proper subset of nodes in another edge, eg, one edge may comprise of $2$ islands, another may comprise of only one of those $2$ islands. Probability that an edge is a target edge can be lower than that for another edge which is its proper subset. For example in Figure 3, consider an edge that is island $4$, and another that is islands $1,4$. The first is a proper subset of the second. Probability that the 1) first is a target edge is $p_{4}(1-p_{1})(1-p_{2})(1-p_{3})=(1-\gamma)(1-\epsilon)1/4\approx 1/4$ 2) second is a target edge is $p_{4}p_{1}(1-p_{2})(1-p_{3})=(1-\gamma)\epsilon 1/4\approx 0.$ The difference arises because $1-p_{1}\approx 1,$ while $p_{1}\approx 0.$
Next, weight of a set $S$ of nodes is not equal to the sum of the probabilities that the nodes in it are infected. Simplest way to see it is to consider $w(S)$ for this example when $S=V$. Since $V$ consists of all the edges, $w(S)$ equals the sum of the probabilities that each edge is a target edge which is $1$ since exactly $1$ edge is a target edge. But note that each node in island $4$ is infected with probability close to $1$. Thus, sum of the probabilities that a node is infected over all nodes in $V$ is at least a number close to $3$ (it is close to $6$ in reality given that there are $2$ moderate islands). Weight of a set $S$ can be 00 even when it contains nodes whose probability of infection is close to $1$. For example consider a $S$ that consists of $2$ nodes in the high island $4$, it consists of $2$ nodes which are infected with probability close to $1$, but it does not consider a single edge because an edge has either all or no node in any given island. Thus $w(S)=0.$
Figure 3: An example with $k=4$ islands. Each edge contains either all nodes in an island or none of it. There are therefore $152^{4}-1=$ 15 edges. Three edges are shown in red. Nodes in an edge can be a proper subset of nodes in another edge, eg, one of the edges marked in red consists of islands $2,3,4$, another consists only island $4$ Nodes in island 1 are infected with $p_{1}=\epsilon\approx 0$, nodes in islands 2 and 3 are infected with probability $p_{2}=p_{3}\approx 1/2$, and nodes in island 4 are infected with high probability $p_{4}=1-\gamma\approx 1$.
Despite the counterintuitive properties above, the following intuitive property holds: $w(V\setminus\{v\})=1-p_{v}.$ This can be argued as follows. Note that when a node $v$ is removed from $S$, for any $S\subseteq V$, $w(S)$ reduces by the sum of probabilities of the edges containing $v$ which were contained in $S$. This sum equals the sum of probabilities of all edges containing $v$ when $S=V$ as then all edges are contained in $S.$ And, the sum of probabilities of all edges containing $v$ equals the probability that the target edge contains $v$ (since only one edge can be the target edge), which in turn equals the probability that $v$ is infected as $v$ is infected if and only if the target edge contains it.
@@ 3.3 Overview and Ideas
In Corollary 4.6, we show that in Stage 1, where each test removes at least $c$ fraction of the mass, the algorithm performs at most $O(H(X))$ tests in expectation which is near-optimal. In Stage 2, where high probability nodes are tested individually, we show in Observation 4.8 that the number of tests depends on the expected number of infections at that point, $\tilde{\mu}$. This expectation is obtained by the posterior probability after performing all the previous tests. We prove bounds on $\tilde{\mu}$ under assumptions on the size of the edges. Specifically, when the size of the edges is concentrated, we get a bound in the form of $O(\mu+H(X))$, which is stated in Theorem 4.9. Another way to bound the number of tests in Stage 2 is to first drop the edges that are much larger than $\mu$, which have a low mass, and then run the algorithm. In this case, we can guarantee that $\tilde{\mu}$ is not much larger than $\mu$, and the guarantee for this approach is shown in Theorem 4.11.
@@ 3.4 The Proposed Algorithm
1 Input: Graph $G=(V,E)$, distribution $\mathcal{D}=\mathcal{D}$ over $E$, $c<1/2$.
@@ 6 If $c\leq w(S)\leq 1-c$, test $V\setminus S$:
7     If the test is positive, remove $E(S)$ from $\mathcal{D}$ and update the posterior distribution following Definition 3.2. Set $W=W\bigcup\limits_{v:q_{v}=0}v$.
8     If the test is negative, remove $E\setminus E(S)$ from $\mathcal{D}$ and update the posterior distribution following Definition 3.2. Set $W=W\bigcup\limits_{v:q_{v}=0}v$.
@@ 10 Else:
15        If the test is positive, remove $E(S)$ from $\mathcal{D}$ and update the posterior distribution following Definition 3.2. Set $W=W\bigcup\limits_{v:q_{v}=0}v$. Go to line 1.
16        If the test is negative, remove $E\setminus E(S)$ from $\mathcal{D}$ and update the posterior distribution following Definition 3.2. Test every node $v\in S$ individually with $10<q_{v}<$ 1, update the posterior after every test and return the nodes with $q_{v}=1$.
@@
The proposed algorithm is presented in Algorithm 1. The algorithm undertakes two broad tasks repeatedly:
@@
    Determine the set of nodes it tests, denoted as $V\setminus S$, it identifies a set $S$ whose complement it tests;
@@
We now describe how these broad tasks are executed. In Stage 1 of the algorithm which corresponds to lines 1-1, the algorithm greedily finds a set $S$ so that $c\leq w(S)\leq 1-c$. This is captured in lines 1 and 1. When the set $S$ is found, $V\setminus S$ is tested in line 1 and the posterior is updated accordingly. Such a test helps rule out edges with total weight of at least $c$ regardless of the outcome of the test. If the algorithm can not find a set $S\setminus v$ so that $c\leq w(S\setminus v)\leq 1-c$ at this stage, it tests $V\setminus S$ in line 1 which has a high probability of being negative. If at line 1, the test result is positive (unexpectedly), the edges that the algorithm rules out constitute of $1-c$ fraction of the mass and the algorithm gets back to line 1. Stage 2 of the algorithm is outlined in line 1, where it tests uncertain high probability nodes individually. Note that at this point, all the nodes outside of $S$ are negative.
@@
We now provide details of the sequence followed in the algorithm. At the start of the algorithm, typically the set of nodes that are known to be negative, $W$, is empty. Thus in line 1, the set $S$ is initialized as $V\setminus W=V.$ Clearly in line 1 we have $w(S)=1$, and $S$ does not satisfy the condition in line 1. The algorithm would now move to line 1 (typically the condition in line 1 would also fail as the target edge will not be clear at this point). In line 1, the algorithm identifies individual nodes which are tested in the loop starting from line 1, and testing these individual nodes rules out edges with total weight at least $c$ regardless of the outcome of the test. After each test, posteriori probabilities are modified and the graph is modified as in Definition 3.2 (this modification happens in lines 1,1 and in line 1). The above sequence is repeated until there does not exist individual nodes such that by removing them $S$ satisfies the condition in line 1. Note that $w(S)=1$ the first time that the algorithm reaches line 1. The tests are all individual while this duration lasts.
@@
Once the duration described in the previous paragraph ends the algorithm moves to line 1. Note that $w(S)=1$ the first time that the algorithm reaches line 1 under the then posterior probabilities. When the algorithm reaches line 1, $S$ consists only of nodes that are positive either with very high probability (such that if any one of those is removed, $w(S)$ falls below $c$ from $1$), or with very low probability (such that $w(S)$ exceeds $1-c$ even when any of those is removed from $S$). In line 1, the low probability nodes are identified one at a time, and once such a low probability node is identified, line 1 determines if now there exists a node whose removal ensures that $w(S)$ satisfies the condition in line 1. If there exists such a node, in line 1 it is tested together with the low probability node identified in line 1. If there does not exist such a node, then the algorithm gets back to line 1 again and removes another low probability node and the same process is repeated. If no node is found in line 1, then all low probability nodes are tested together in line 1. Thus, tests are no longer individual once the algorithm reaches line 1. When a group of nodes is tested, either in line 1 or line 1, the weights are modified as in Definition 3.2. If the test in line 1 is negative, then all tested nodes are removed from $V$, and the remaining nodes are all positive with high probability and are tested individually in line 1. If the test is positive, then only edges consisting of at least $1$ node in the set are retained, and the algorithm returns to line 1 and the sequence is repeated.